8.10 kJ of heat energy is added to 15.0 g of ice at −25.0 °C. Describe the final state of the system and its temperature.
What state of matter is it? (solid, liquid, gas)
what is the final temperature (C)
1
Expert's answer
2018-01-22T14:01:08-0500
C (ice) = 2.05 J / g · K*
The formula for calculation: Q = cmDT
So DT = Q / cm = 8,100 J / 2.05 J / g · K · 15.0 g = 263.4 K
If the temperature of ice is -25.0 °C the heat added causes the increase of temperature DT = 263.4 K. This is about 238.4 °C which is a water vapour.